OpenStudy (anonymous):

The distance traveled, in feet, of a ball dropped from a tall building is modeled by the equation d(t) = 16t2 where d equals the distance traveled at time t seconds and t equals the time in seconds. What does the average rate of change of d(t) from t = 2 to t = 5 represent?

OpenStudy (anonymous):

The ball travels an average distance of 112 feet from 2 seconds to 5 seconds. The ball falls down with an average speed of 48 feet per second from 2 seconds to 5 seconds. The ball falls down with an average speed of 112 feet per second from 2 seconds to 5 seconds. The ball travels an average distance of 48 feet from 2 seconds to 5 seconds.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@luckycoins888 @TheSmartOne @mathmate How come I can't get anyone to help me with this?

OpenStudy (mathmate):

@CountryGurl15 have you worked with calculus yet?

OpenStudy (anonymous):

Um not to my knowledge =)

OpenStudy (mathmate):

So the word derivative, or dy/dx has not appeared in your math or physics class?

OpenStudy (mathmate):

Is this physics or math question? Because in physics, there are standard equations for velocity, but in math, you're not supposed to know them.

OpenStudy (mathmate):

@CountryGurl15 Have you looked at the total and average distance travelled between 2 and 5 seconds, and compare with the given answers?

OpenStudy (anonymous):

Um, no it has not =)

OpenStudy (anonymous):

this is an Algebra 1 question

OpenStudy (mathmate):

Well, d(t)=16t^2 gives you the distance in feet at time t (seconds). So d(2)=16(2^2)=64 is the distance at t=2 seconds, and d(5)=16(5^2)=400 is the distance at t=5 seconds. So the ball travelled a distance of (400-64)=336 feet in 3 seconds. That's and average of 336 feet /3 seconds = 112 feet/second That's all the math involved. Now the work is for you to choose the right choice. Hint: units are very important in science and math.

OpenStudy (anonymous):

Thank you so much for the help! It really helped me =)